OpenStudy (anonymous):

A question #9 r and s are two different real number and satisfied (r/s)+[(r+10s)/(s+10r)]=2 find r/s

OpenStudy (anonymous):

might be an easy one for you guy :P

OpenStudy (mr.math):

Let me just write it using latex: \[\frac{r}{s}+\frac{r+10s}{s+10r}=2\]

OpenStudy (anonymous):

hmm \[2 - r/s = \frac{r/s + 10}{10r/s +1}\]

OpenStudy (anonymous):

r/s = x

OpenStudy (anonymous):

but i am not getting get good roots

OpenStudy (anonymous):

can we use latex here??

OpenStudy (anonymous):

\[2 - x = \frac{x + 10}{10x +1}\]

OpenStudy (anonymous):

oh 4/5and1

OpenStudy (anonymous):

i mean i never know we can use it here

OpenStudy (anonymous):

so wats ur ans??

OpenStudy (anonymous):

sorry these are good roots i made a mistake earlier

OpenStudy (anonymous):

4/5 and 1

OpenStudy (anonymous):

read the problem carefully LOL

OpenStudy (lollylau):

...

OpenStudy (anonymous):

and i think i read the problem carefully

OpenStudy (anonymous):

It require 2nd equation for solving two variables...

OpenStudy (lollylau):

write y in terms of x -.-

OpenStudy (anonymous):

I got 4/5 to r=4 and s=5.

OpenStudy (anonymous):

ishaan & foolfm, both of you are doing in the correct way. and your answer is nearly correct because r/s=4/5 and 1 truly but from the problem "r and s are two different real number" so the answer should go to 4/5 or 0.8 only LOL

OpenStudy (anonymous):

Yes, I was going to post the same thing, when my internet got disconnected. Only 4/5 is the correct answer to this problem.